### v2

# “I know myself better, but not really greatly”: How Well Can LLMs Detect and Explain LLM-Generated Texts?

Abstract
Distinguishing between human- and LLM-generated texts is crucial given the risks associated with misuse of LLMs. This paper investigates detection and explanation capabilities of current LLMs across two settings: binary (human vs. LLM-generated) and ternary classification (including an “undecided” class). We evaluate 6 close- and open-source LLMs of varying sizes and find that self-detection (LLMs identifying their own outputs) consistently outperforms cross-detection (identifying outputs from other LLMs), though both remain suboptimal. Introducing a ternary classification framework improves both detection accuracy and explanation quality across all models. Through comprehensive quantitative and qualitative analyses using our human-annotated dataset, we identify key explanation failures, primarily reliance on inaccurate features, hallucinations, and flawed reasoning. Our findings underscore the limitations of current LLMs in self-detection and self-explanation, highlighting the need for further research to address overfitting and enhance generalizability.
“I know myself better, but not really greatly”: How Well Can LLMs Detect and Explain LLM-Generated Texts?
[…]
1 Introduction
The rise of large language models (LLMs) has brought remarkable advancements in natural language processing (NLP) tasks matarazzo2025survey, including text generation. Models such as GPT-4o OpenAI2024GTP4O, LLaMA Touvron2023LLaMA, and Qwen ali2024qwen2 have blurred the boundaries between LLM-generated (LGTs) and human-generated texts (HGTs), posing new challenges in distinguishing between the two. While these capabilities of LLMs open new possibilities, they also bring concerns in areas such as misinformation, academic dishonesty, and automated content moderation hu2025unveiling. As a result, detecting LGTs has become an increasingly important research area dugan2024raid; lee2023do; bhattacharjee2024fighting.
Prior research has mainly focused on developing classifiers to distinguish HGTs and LGTs, including open-source detectors Hans2024Binoculars and online close-source detection systems gptzero_website. However, most detection systems have been limited to binary classification, which has several inherent issues. Recently, some works lee2024llm have attempted ternary classification by introducing a “mixed” category, which represents texts originating from mixed sources. However, this approach does not fundamentally resolve the issue. We further adopt the definition of an “Undecided” category based on other studies ji2024detecting and conduct ternary classification experiments for different LLMs, as certain texts are inherently indistinguishable between LGTs and HGTs. Furthermore, many studies treat the detection task as a black box, offering little insight into the decision-making process. Explainability, a critical aspect of trustworthy AI, has received less attention, but it is essential for building systems that users can trust weng2024understanding; zhou2024humanizing. This paper presents an analysis of current LLMs in detecting LGTs and HGTs, with a particular emphasis on evaluating and improving the clarity of the explanations provided by LLM-based detectors. By investigating how LLMs make predictions and offer explanations for their decisions, we aim to enhance their transparency and provide deeper insights into their reasoning processes.
This paper explores the explainability of LLM-based detectors, addressing two central questions: (1) How accurately can current LLM detectors identify origins of texts, and (2) How reliable are their explanations? Our study highlights that in a ternary setting compared to traditional binary classification, the average detection performance improves by 5.6%, which demonstrates the necessity of ternary rather than binary setting to detect HGTs and LGTs. We further discovered that explanations are often flawed even when binary predictions are correct. Based on our comprehensive human-annotators’ feedback, we summarize three common issues with explanations: reliance on inaccurate features (e.g., vague or irrelevant characteristics), hallucinations (e.g., non-existent or contradictory features), and incorrect reasoning (e.g., logical errors in attributing text origin). These explanation errors are quantified and categorized, with their distributions analyzed across different LLMs. Consequently, the proportion of explanation errors decreases by 13.3% when we switch to ternary classification setting, which further supports the necessity of ternary classification for LGTs detection.
We evaluated 6 state-of-the-art (SOTA) LLM-based detectors, such as GPT-4o, GPT-4o mini, LLaMA3.3-70B, LLaMA3.3-7B, Qwen2-72B, and Qwen2-7B, on our created dataset comprising LGTs and HGTs. Moreover, our human annotators provided feedback based on correctness of predictions and explanations for this benchmark. Our results show that GPT-4o achieved the highest detection accuracy. In addition, LLMs performed better in self-detection than cross-detection, and ternary classification outperformed binary classification. Finally, explanation quality also improved under ternary setting, with fewer hallucinations and incorrect reasoning observed.
The main contributions of this work are:
  • Comprehensive Evaluation of Detection and Explanation: We systematically assess current LLMs’ ability to detect and explain human- and LLM-generated texts using both binary and ternary classification tasks, demonstrating the advantages of ternary classification for both detection accuracy and explanation quality.
  • Human-Annotated Dataset: We present a new human-annotated dataset of LLM- and human-generated texts, enabling evaluation of LLM explanations and improved detector training.
  • Analysis of Explanation Errors: We quantitatively and qualitatively characterize key explanation failures, reliance on inaccurate features, hallucinations, and flawed reasoning, offering insights for LLM detection and self-explanation.
Datasets Models LLM Detectors
GPT-4o GPT-4o mini LLaMA3.3-70B LLaMA3.3-8B Qwen2-72B Qwen2-7B
GPT-4o 71.39 59.38 57.31 48.41 64.14 59.76
GPT-4o mini 65.71 61.03 53.75 51.73 67.27 60.09
LLaMA3.3-70B 67.26 60.92 68.10 53.65 58.96 51.57
LLaMA3.3-8B 60.74 55.77 62.29 59.09 59.87 49.88
Qwen2-72B 62.66 61.92 57.79 49.20 68.15 61.36
Qwen2-7B 62.45 59.06 59.12 48.57 65.24 63.44
Average 65.03 59.68 59.73 51.78 63.94 57.68
Table 1: F1 scores of LLM-based detectors in binary classification. The first column indicates different LLMs used for text generation, and the first row indicates different LLMs acting as detectors. The highest column-wise F1 score for each LLM detector to classify LGTs and HGTs across six datasets is highlighted in blue. The highest row-wise F1 score for each LLM-generated text dataset across different LLM detectors is marked in blue.
2 Related Work
LGT and HGT Detection.
Past efforts to identify LGTs often relied on binary classification systems that distinguish HGTs from LGTs using surface-level features. While these methods were initially effective, they are prone to errors when encountering adversarial attacks or domain shifts, which limit their overall robustness bhattacharjee2024fighting; dugan2024raid. To address these limitations, researchers have explored strategies that integrate external knowledge, such as combining internal and external factual structures, to boost detection against diverse content and styles ideate2024. Recent studies also highlight the promise of using LLMs themselves for text detection: approaches like self-detection and mutual detection can outperform traditional classifiers, as illustrated by GPT-4’s success in tasks like plagiarism detection plagbench2024. Notably, smaller models sometimes excel in zero-shot scenarios, offering adaptable solutions across varying architectures smallerLLM2024. Furthermore, lee2023do demonstrated that LLMs can reliably identify their own outputs, providing a more nuanced framework for content verification. Despite these advances, the continuing challenges of domain adaptation and adversarial resistance underscore the need for more versatile and robust detection systems.
Explainability in Detection Models. Recent work on LGT detection has focused on improving explainability. zhou2024humanizing proposed to incorporate factual consistency into detection models to enhance their interpretability, while weng2024understanding explored mixed-initiative approaches that combine human expertise with automated models for better detection. These studies have made significant contributions to the field; however, they either depend heavily on expert input weng2024understanding or lack integration of explanation generation within the model itself zhou2024humanizing. Our approach, in contrast, enables LLMs to autonomously generate both predictions and detailed explanations, making it a more scalable and transparent solution for detecting machine-generated content.
Ternary Classification. Traditional binary classification methods face limitations when texts exhibit ambiguous characteristics. Introducing an “Undecided” category addresses this by capturing three distinct scenarios (see Appendix LABEL:app:undecided_examples for concrete illustrations): (1) Mixed texts co-authored by humans and LLMs, where stylistic blending creates classification challenges; (2) Inherently ambiguous texts that could plausibly originate from either source despite single authorship; and (3) Fragile indicators, where subtle distinguishing features exist but lack robustness against behavioral evolution of either LLMs or human writers. This approach advances beyond previous methods that primarily addressed mixed texts lee2024llm. The complexity is evidenced by Turing tests showing human difficulty in binary classification turingtest2024, and by studies demonstrating detector limitations with evolving writing patterns coling2024. The ternary framework improves both accuracy and explainability, particularly for these edge cases.
3 LLM-based Binary Classification on LGTs and HGTs
3.1 Experimental Design
We selected six SOTA LLMs for text generation and subsequent detection: GPT-4o, GPT-4o mini hurst2024gpt, Qwen2-72B, Qwen2-7B yang2024qwen2, LLaMA3.3-70B, and LLaMA3.3-8B dubey2024llama. These LLMs were chosen for two main reasons. First, they represent the latest advancements in LLM development, demonstrating strong generation and detection capabilities. Second, selection spans different series and model sizes, enabling a comparative analysis of performance across architectures and scales.
To construct the dataset, we first selected 1,000 HGTs from publicly available M4GT-Bench dataset wang2024m4gt, ensuring a diverse range of topics, styles, and formats. Based on these selected HGTs, we designed 1,000 prompts that align with themes, structure, and style of the HGTs. Each LLM subsequently generated a corresponding response for each of these prompts. Together, these LGTs and HGTs formed the benchmark used in this study. For each text, the LLMs were tasked to determine its source (LGTs or HGTs) and provide an explanation, as illustrated in Table 2.
Prompt: Please determine whether the following text is generated by large language models or by a human, and provide a clear judgment. Additionally, please offer a detailed explanation for your decision. Please structure your answer in JSON format as follows: {“answer”: , “explanation”: }.
Table 2: A prompt for LLMs to determine text origin and provide an explanation under a binary setting.
Datasets Models LLM Detectors
GPT-4o GPT-4o mini LLaMA3.3-70B LLaMA3.3-8B Qwen2-72B Qwen2-7B
GPT-4o 73.48 / 67.04 58.41 / 54.17 57.65 / 54.17 48.29 / 51.32 64.11 / 59.60 59.57 / 62.13
GPT-4o mini 63.72 / 60.95 63.43 / 60.15 53.85 / 52.46 50.01 / 47.75 66.91 / 61.08 60.13 / 57.28
LLaMA3.3-70B 68.13 / 63.96 62.12 / 60.32 68.33 / 64.47 53.13 / 51.78 58.16 / 59.22 52.41 / 48.82
LLaMA3.3-8B 58.97 / 61.11 56.19 / 55.98 63.24 / 63.72 58.73 / 56.29 59.83 / 59.17 49.66 / 48.97
Qwen2-72B 62.70 / 61.09 62.91 / 62.84 58.71 / 56.99 49.12 / 47.26 70.47 / 67.98 61.24 / 58.18
Qwen2-7B 63.78 / 61.54 58.11 / 57.84 60.15 / 58.60 48.72 / 49.17 65.44 / 63.58 63.83 / 61.91
Table 3: F1 scores of LLM-based detectors on human-annotated texts for binary classification. Each dataset contains 100 LGTs and 100 HGTs with human-annotated explanations. Each cell indicates classification/explanation F1, where the highest column-wise F1 of each LLM detector for binary classification and explanations across different generated texts are highlighted with blue and red, respectively. In addition, the highest row-wise F1 among different LLM detectors for each LLM-generated text datasets are indicated with blue and red in bold, respectively.
Manual Annotation.
To assess LLMs’ ability to explain text origins and identify distinguishing features, 3 co-authors, who are undergraduate computer science students, manually evaluated correctness of LLM-generated explanations. They determined accuracy of each explanation. From 7 datasets (6 SOTA LLMs + Human), 100 texts with corresponding explanations per dataset were randomly selected for human evaluation. All annotators assessed explanations provided by each model, which achieved a Fleiss’ kappa Fleiss1971kappa of 0.8387, indicating near-complete agreement. Annotation guidelines are detailed in Appendix LABEL:appendix:annotation_guidelines.
Evaluation Metrics.
For evaluating the classification performance of the LLMs, the primary metric we used is the F1 score. To assess the quality of explanations, human evaluators reviewed the LLM-generated explanations and classified them as correct or incorrect. The F1 score was also used as the evaluation metric for explanation quality.
3.2 Binary Classification Results
We evaluated the performance of six LLMs across six datasets, as detailed in Table 1, which systematically compares the detection capabilities of various LLMs for both LGTs and HGTs. The results demonstrate that GPT-4o achieves the best average detection performance across all datasets, showing relatively strong generalization capabilities. Larger parameter models generally exhibit significantly better detection performance than smaller ones, which suggests that these models are not merely making random guesses but are effectively identifying distinctive textual features.
The F1 scores in Table 1’s diagonal direction show that LLMs within same series consistently detect their own outputs more effectively than those from other LLM families. For example, LLaMA3.3 70B achieves the highest F1 score in its generated dataset, which indicates a heightened sensitivity to its own text distribution compared to other LLMs. However, this specialization reduces cross-detection performance, as seen in Qwen2-7B’s lower F1 on LLaMA-generated texts. While larger LLMs generally achieve better detection across different LLMs, such as GPT-4o, GPT-4o mini, LLaMA3.3-70B and Qwen2-72B, their outputs are also more difficulty to distinguish by smaller LLMs, such as LLaMA3.3-8B and Qwen2-7B.
Additionally, based on the human annotations of sampled 100 LGTs and 100 HGTs with explanations from each dataset, we observed that the detection and explanation results across different LLMs are not entirely consistent, as shown in Table 3. We noted that in some cases, the F1 score for explanations was higher than that for classification. This is because, in these cases, the explanation correctly identified the reasoning for attribution, but the final classification was incorrect. For instance, the difference in F1 scores between explanation and classification was particularly noticeable for LLaMA3.3-8B and Qwen2-7B, suggesting that these models struggle to truly comprehend the textual features necessary for correctly determining the origin of generated texts, which results in lower detection performance.
As shown in Table 4, analysis of the annotators’ results revealed that models are generally more accurate in attributing HGTs compared to LGTs. For example, while GPT-4o demonstrates higher accuracy (78 out of 100) in classifying HGTs, the false explanations account for more than 47%.
Model MGTs HGTs
TC TE FE FC TE FE TC TE FE FC TE FE
GPT-4o 64 ${51}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:79.7\%}}$ ${13}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:20.3\%}}$ 36 ${8}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:22.2\%}}$ ${28}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:77.8\%}}$ 78 ${41}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:52.6\%}}$ ${37}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:47.4\%}}$ 22 ${2}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:9.1\%}}$ ${20}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:90.9\%}}$
LLaMA3.3-70B 56 ${35}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:62.5\%}}$ ${21}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:37.5\%}}$ 44 ${10}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:22.7\%}}$ ${34}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:77.3\%}}$ 60 ${37}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:61.7\%}}$ ${23}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:38.3\%}}$ 40 ${7}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:17.5\%}}$ ${33}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:82.5\%}}$
Qwen2-72B 60 ${36}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:60.0\%}}$ ${24}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:40.0\%}}$ 40 ${7}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:17.5\%}}$ ${33}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:82.5\%}}$ 69 ${50}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:72.5\%}}$ ${19}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:27.5\%}}$ 31 ${7}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:22.6\%}}$ ${24}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}:77.4\%}}$
Table 4: Performance of LLMs on LLM-generated and human-generated texts for ternary classification and explanation tasks. It includes results for classification and explanation tasks, where TC represents true classification, FC represents false classification, TE represents true explanation, and FE represents false explanation. Note: TC=TE+FE and FC=TE+FE.
Datasets Models LLM Detectors
GPT-4o GPT-4o mini LLaMA3.3-70B LLaMA3.3-8B Qwen2-72B Qwen2-7B
GPT-4o 79.73/72.04 64.62/61.87 62.19/59.04 58.06/57.78 71.62/68.86 63.81/62.72
GPT-4o mini 70.11/68.75 67.39/ 65.18 58.88/52.95 54.43/51.16 69.65/65.95 65.15/62.60
LLaMA3.3-70B 74.41/ 75.26 65.16/64.75 72.11/ 71.83 57.05/57.34 64.94/62.44 56.46/55.32
LLaMA3.3-8B 71.99/70.80 60.18/61.10 64.82/63.93 63.96/ 62.85 63.12/60.52 54.08/53.01
Qwen2-72B 67.28/66.74 65.12/64.73 61.81/61.74 53.24/52.87 76.05/ 75.56 65.26/64.72
Qwen2-7B 68.91/67.42 60.15/59.31 62.06/61.57 52.41/52.30 70.30/68.44 66.61/ 65.17
Table 5: F1 scores of LLM-based detectors on the ternary classification of LGTs and HGTs. The highest column-wise and row-wise F1 scores are highlighted and marked following the same scheme as in Table 3.
Prompt: Please classify the following text into one of three categories based on its source: LLM-generated, human-generated, or undecided. The “Undecided” category refers to texts that exhibit characteristics of both LLM-generated and human-generated content, making it impossible and inappropriate to distinguish between the two. Provide a clear classification and a detailed explanation for your decision. Structure your answer in JSON format as follows: {“classification”: , “explanation”: }.
Table 6: A prompt for LLMs to determine text origin and provide an explanation under a ternary setting.
4 LLM-based Ternary Classification on LGTs and HGTs
4.1 Experimental Setup
Using the same benchmark in § 3, we prompted the LLMs for ternary classification and the prompt template is demonstrated in Table 6. The ground truth for the ternary classification was determined based on annotators’ votes, where the three annotators were aware of the text’s origin (LLMs or human) and were asked to distinguish between the ground truth and the “Undecided” category. This allowed for the evaluation of both the LLM’s classification results and the explanations provided by the LLMs. The Fleiss’ kappa Fleiss1971kappa for the ternary classification annotations among the three annotators was calculated as 0.7629, which indicates substantial agreement.
4.2 Ternary Classification Results
Table 5 presents the F1 scores of LLMs in the ternary classification setting. Comparing it with Table 3, we observe that introducing the “Undecided” category leads to overall performance improvements across both classification and explanation tasks. Specifically, GPT-4o exhibits the most notable gains, improving from 73.48/67.04 to 79.73/72.04, indicating that a finer-grained classification allows stronger models to better capture nuanced differences between LGTs and HGTs.
Moreover, Fig LABEL:fig:ternary_confusion_matrices reveals how different models distribute predictions across three categories. GPT-4o demonstrates a more balanced distribution, with relatively lower misclassification rates for both HGTs and LGTs. In contrast, LLaMA3-70B shows a stronger tendency to label texts as “human-generated”, leading to a higher false positive rate. Meanwhile, Qwen2-72B exhibits a more cautious classification approach, assigning a larger proportion of texts to “Undecided” category, particularly for LGTs.